has an average of 0.8 flaws per 4 meter length of wire. 4 4. A roll of plastic coted wire Suppose a quality engineer will sample 3 meter length of wire from a roll of 220 meters in length. If no flaws are found on the sample the engineer will accept the entire roll. What is the probability that the roll will be rejected? (3 pts)
Answer #1

Answer:

for 4meters 0.8flaws given...so that for 8meters it will be 1.6

Total number of flas in total length of the wire = 220 * 1.6 /8 = 44

sampling taken is 8m,

Probability that one flaw to be in 212m = 212/220

for 44 flaws it will become (212/220) ^44 = 0.2

probability that the wire to be rejected = 1 - 0.2 = 0.8

probability that the roll will be rejected because one flaw as found in the sample

= (8/220) * (212/220) ^ 43

=0.008